OpenStudy (anonymous):

If f(x) = log2 (x + 4), what is f−1(3)? 0 2 4 8

hartnn (hartnn):

log2 (x+4) = 3 find x from here :)

OpenStudy (anonymous):

is it 0?

OpenStudy (jdoe0001):

\(\bf f(x)={\color{blue}{ y}}=log_2({\color{red}{ x}}+4)\qquad inverse\implies {\color{red}{ x}}=log_2({\color{blue}{ y}}+4)\\ \quad \\ \quad \\ \textit{log cancellation rule of }{\color{red}{ a}}^{log_{\color{red}{ a}}\square }=\square \qquad thus\\ \quad \\ x=log_2(y+4)\implies {\color{red}{ 2}}^x={\color{red}{ 2}}^{log_{\color{red}{ 2}}(y+4)}\implies {\color{red}{ 2}}^x=y+4\)

OpenStudy (jdoe0001):

solve for "y" and plug in the "3"

OpenStudy (anonymous):

is it 4 then? @jdoe0001

OpenStudy (jdoe0001):

well... what did you get for the inverse?

OpenStudy (anonymous):

I don't know how to do this, I'm just trying to narrow it down.

OpenStudy (jdoe0001):

well, that's the idea.... that you understand it, thus the exercise

OpenStudy (mertsj):

In the inverse function, all the ordered pairs are reversed. If 3 is an x value of the inverse function, that means it was a y value in the function.

OpenStudy (mertsj):

That is what hartnn was trying to show you. Take the function and replace y with 3

OpenStudy (mertsj):

\[\log_{2}(x+4)=3 \]

OpenStudy (jdoe0001):

yeap

OpenStudy (mertsj):

Now put it into its exponential form

OpenStudy (mertsj):

\[2^3=x+4\]

OpenStudy (jdoe0001):

@tornjeansxo have youj covered inverse functions yet?

OpenStudy (anonymous):

I learned all of this but forgot. It was a few months ago so :/

OpenStudy (mertsj):

So we see that x = 4 and (4,3) is an ordered pair of the original function. Therefore (3,4) is an ordered pair of the inverse function and: \[f ^{-1}(3)=4\]

OpenStudy (anonymous):

thank you, guys :)

OpenStudy (mertsj):

yw